# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v1

@@ IV-A Overview of Methods
Figure 5: Architecture and workflow of the Direct NVMe Engine for optimized SSD I/O.
IV-B Adaptive Buffer Pool
To mitigate system memory fragmentation, a precise understanding of buffer requirements is essential. These requirements are primarily driven by data transfer demands between SSDs and GPUs during layer-wise forward and backward propagation and commonly employ asynchronous prefetching. Due to pre-fetching operations that will cause N number of blocks concurrently in flight, the required buffer capacity is for N transformer blocks. In each transformer blocks, since smaller tensors can remain permanently in CPU memory, the primary tensors requiring transfer include the feedforward layer weights, specifically the up, down, and gate projections, and the attention layer weights, namely the Q, K, V, and O projections. Additionally, dedicated buffers are required for the embedding layer weights and the final linear layer weights, often referred to as the LM head. Thus, the overall buffering strategy must accommodate weights for one embedding layer, N transformer blocks each containing multiple weight tensors, and one LM head.
Compared to the original SSD offloading system, which suffers from internal fragmentation inherent in the monolithic buffer pool approach, our proposed adaptive memory buffer pool architecture addresses this limitation by implementing distinct pools for different tensor shapes. Given that embedding and LM head weights share identical dimensions (vocabulary size $\times$ hidden size), and feedforward projections maintain consistent dimensions (intermediate size $\times$ hidden size), we establish four specialized pools: embedding-related weights, feedforward projections, K and V projections (commonly identical in models employing grouped-query attention), and Q and O projections. These groups maintain counts of 2, 3N, 2N, and 2N, respectively, effectively eliminating internal fragmentation.
This approach is effective not only for single-GPU setups, in multi-GPU configurations employing parameter partitioning, each process fetches only its assigned shard of a tensor. Consequently, the individual buffer sizes required per process are reduced proportionally to the number of partitions. However, from a system wide perspective, the aggregate buffer capacity remains consistent with single GPU configurations. Therefore, the proposed adaptive buffer pool optimization maintains its effectiveness across various distributed GPU setups.
As illustrated in Figure 5, this adaptive approach utilizes specialized subgroups for different tensor sizes. This method significantly mitigates the internal fragmentation observed in the monolithic buffer pool characteristic of original SSD offloading systems.
IV-C Zero Overhead Pinned Memory Allocation
@@
As noted in Section III-B, PyTorch’s default power-of-two alignment for pinned memory allocations causes significant memory waste, especially with large static buffers used in SSD offloading. To avoid this overhead, the standard PyTorch caching allocator is bypassed for these buffers. Instead, memory is allocated directly using system calls and CUDA runtime APIs. The required host memory is first allocated with standard memory functions, aligned only to the minimal extent required by the system. This avoids the excessive rounding imposed by power-of-two policies. The allocated region is then registered as pinned using the CUDA API cudaHostRegister, making it accessible for efficient GPU DMA transfers. This direct allocation and registration strategy ensures that pinned memory buffers use only the space needed for the data plus minimal alignment. It eliminates the internal fragmentation caused by power-of-two rounding and substantially reduces system memory usage for large allocations such as the gradient partition flat buffer and the adaptive buffer pool.
IV-D Fused Overflow Check
As mentioned in the motivation section, when the SSD offloading system checks for gradient overflows, it involves multiple operations such as isnan(), isinf(), and others. Each of these operations in PyTorch produces several intermediate values, resulting in a peak memory overhead of approximately 2.25 times the size of the flat buffer of the gradient partition, which corresponds to the model size. Performing multiple steps for a single task is inefficient. By fusing these operations into a single operation, we can reduce memory usage and improve performance.
The fused overflow check logic is illustrated in Figure 2(b). In the SSD offloading system, the flat buffer of the gradient partition is typically stored in fp32 to support gradient accumulation. While fp32 helps prevent overflow, the gradients computed during backpropagation remain in fp16, meaning overflow is still possible and must be checked in the fp32 tensors.
According to the IEEE 754 standard [29], NaN values have all exponent bits set to 1 and a non-zero mantissa, whereas positive and negative infinity have exponent bits set to 1 and a zero mantissa. Leveraging this property, we can efficiently detect whether any value has an exponent field of all 1s using bit-wise operations. If this condition is met, the value is considered to have overflowed; otherwise, it has not. This eliminates the need for intermediate values. Furthermore, as this operation is dependency-free, it can be parallelized to maximize performance. This fused overflow check achieves an average efficiency of approximately 97% across model sizes ranging from 1 billion to 32 billion parameters, and resolves the bottleneck issue that arises during overflow checks at large scale.
@@ IV-E Direct NVMe Engine
The primary goal of the Direct NVMe Engine is to bypass the Linux filesystem overhead by using raw I/O to directly send asynchronous I/O (AIO) requests to the device driver via iouring or libaio. The overview of architecture and its workflow is shown in Figure 5.
First, the SSD offloading system issues a tensor AIO request. The request context contains a unique tensor name (as a key), a data pointer, and an offset aligned to the tensor size. Next, the Direct NVMe Engine distributes this request among multiple thread workers, dividing the data into equal portions for balanced performance. Each thread consults the Tensor Location Dictionary to determine which SSD device and logical block address (LBA) to use.
If the Tensor Location Dictionary has not recorded the tensor yet, it consults the Location Allocator for SSD space. This step ensures that data are horizontally partitioned across SSDs, effectively implementing striping in place of Software RAID 0. The Shared Device Information structure, implemented via shared memory, guarantees that multiple processes do not allocate overlapping SSD regions.
After each thread obtains its data pointer, offset, LBA, and raw SSD device path, it sends a raw AIO request through libaio. By applying load balancing, stripping, and multi-threading, the Direct NVMe Engine reduces latency and achieves higher throughput compared to approaches that rely on filesystem-based AIO.
@@ V Memory Efficiency and Performance Analysis
MemAscend significantly reduces system memory usage compared to prior SSD offloading systems without sacrificing performance. Lower system memory consumption is key to reducing the cost of training tasks. Conserving system memory offers several key advantages. First, with equivalent hardware, this approach enables training models with more trainable parameters, thus achieving the same training setup at a lower cost. Second, this method supports longer context lengths. Extended context lengths necessitate larger activation checkpointing values. By offloading this component from GPU memory to system memory, the available system memory enables these longer context lengths. Finally, this approach improves training performance. While a primary drawback of SSD offloading is the reduced training throughput stemming from PCIe transfer overhead, offloading also enables larger batch sizes. This boosts GPU efficiency by processing more data per batch. The following sections analyze memory efficiency in detail. Figure 8 shows the memory breakdown across system components during training of the Qwen2.5-7B model [30]. MemAscend shows notable improvements in the key areas of memory usage. The discussion then turns to how reduced memory requirements enable longer context lengths and larger batch sizes, both of which can contribute to higher throughput.
@@ V-A Memory Efficiency
MemAscend first introduces an Adaptive Buffer Pool to reduce memory fragmentation. For the Qwen2.5-7B model in fp16, this approach requires about 2.46GiB of memory for buffer allocations. In contrast, the standard SSD offloading method reserves a fixed 9.14GiB. This change cuts memory use in the parameter buffer pool by 73%. To eliminate peak memory spikes from overflow checking, MemAscend implements a Fused Overflow Check kernel. Conventional checks rely on intermediate buffers, adding a 1.25$\times$ overhead and consuming 35.46GiB. The fused kernel maintains numerical stability without this overhead, reducing usage to 0GiB. SSD offloading also depends on pinned memory to move data between CPU and GPU. Traditional allocation uses power-of-two alignment, introducing significant waste. MemAscend applies a zero-overhead strategy that allocates only the required memory. This reduces pinned memory overhead from 24.90GiB to 1.63GiB, a 93.5% decrease. Some components remain constant across methods, including the Gradient Flat Buffer and small system allocations like optimizer state buffers and the swap-out buffer. These account for stable memory usage of roughly 28.37GiB and 11.17GiB, respectively. MemAscend brings total peak system memory usage down from 109.04GiB under ZeRO-Infinity to 43.64GiB, a 60% reduction. This allows for training larger models and handling longer input sequences within the same memory budget.
V-B Enable Larger Context Length
In the SSD offloading scenario, the key constraint in expanding the context length or batch size is the activation checkpoint value, which scales linearly with both parameters. To alleviate this GPU memory bottleneck, a pinned memory buffer on the CPU can be utilized to swap activation checkpoint values from GPU memory to system memory, potentially enabling longer context lengths. However, this shifts the bottleneck from GPU memory to system memory.
The total size of activation checkpointed values in system memory is calculated using the following formula, where we define $N_{g}$ as the number of GPUs, $B$ as the batch size, $C$ as the context length, $L$ as the number of layers, $H$ as the hidden size, $F_{16}$ as the bytes per fp16, and $P_{m}$ as the pinned memory overhead. Here, the number of layers $L$ accounts for the activation checkpointing enabled in each transformer layer.
@@ V-B Enable Larger Context Length
Figure 8 shows that the usage of system memory increases with increasing context length when running Qwen2.5-7B on two GPUs. Our MemAscend method scales linearly with context length, exhibiting minimal pinned memory overhead, meaning the increase is almost entirely due to context length scaling. In contrast, the original SSD offloading system scales more rapidly and begins with higher memory usage. This disparity arises because its pinned overhead aligns to the nearest power-of-two value, resulting in faster scaling and instances where different context lengths yield identical memory usage due to this alignment. Consequently, our method significantly reduces system memory consumption and slows the scaling rate with increasing context length, offering greater flexibility to extend context length without exhausting system memory. For example, with a 128 GiB system memory limit, the original SSD offloading system supports a context length of 16,384, while MemAscend scales up to 131,072 under the same configuration.
V-C Enable Higher Performance
Our MemAscend not only supports a larger trainable context length, but also unlocks the potential for peak performance by enabling a larger batch size. A larger batch size allows the GPU to process more data within the same data transfer volume. In an SSD offloading scenario, performance degradation is primarily due to additional data transfers. By optimizing the ratio of tensor data computation to data transfer processing more data under the same transfer load, we can mitigate the drawbacks of SSD offloading. Specifically, a larger batch size means offloading and retrieving the same volume of data while computing more data, thus enhancing performance potential. As noted in the previous section, the size of the activation checkpoint values in the system memory follows Equation 1. Thus, with greater system memory availability, we can offload larger activation checkpointed values from the GPU, enhancing batch size potential.
In Figure 8, we compare the original SSD offloading system with our MemAscend method during Qwen2.5-7B training on two GPUs. Initially, we observe memory usage and scaling rates consistent with the previous section, as increasing batch size or context length equally impacts the size of activation checkpointed values. And our method exhibits significantly lower system memory requirements and a slower scaling rate across batch sizes. This reduced demand provides MemAscend with the flexibility to increase batch size without risking system memory exhaustion.
Next, we observe that throughput increases nearly linearly with batch size. This confirms that a larger batch size boosts overall performance, as we offload the same data volume while processing more computations. In summary, MemAscend improves performance potential by enabling larger batch sizes with increased data offloading to system memory. For instance, with a 128 GiB system memory limit, the original SSD offloading system supports a batch size of 4, while MemAscend achieves 32 or greater in the same configuration, delivering substantially greater performance.
@@ V-C Enable Higher Performance
VI Experimental Evaluation
VI-A Setup
All experiments were carried out on two servers. Detailed hardware specifications are provided in Table II. MemAscend is evaluated using mixed-precision training of large language models with SSD offloading. The workloads vary by model architecture, context length, and batch size. Models include Llama3.1-8B and Qwen2.5, with parameter sizes of 7 billion, 14 billion and 32 billion. All models use the Hugging Face Transformers library (v4.46.3). The context lengths range from 4,096 to 131,072 tokens. Batch sizes vary across different experiments. The main evaluation metric is peak system memory usage during training. To assess overall performance, we also measure training throughput in tokens per second.
To evaluate the effectiveness of MemAscend, we compare it with ZeRO-Infinity. Both systems share a common setup to ensure fairness. Static memory components, including model parameters and optimizer states, are stored entirely on the SSD. To reduce GPU memory usage, we enable gradient checkpointing, Flash-Attention, and the Liger kernel. For experiments that require offloading gradient checkpoint activations, we use a modified version of Unsloth, adapted from its original single-GPU design.
@@ VI-B1 Assessment of Individual Methods
As each model varies in weight tensor size due to differences in vocabulary size, hidden size, and other factors, this experimental section presents the memory usage of the Parameter Buffer Pool system across various models. The aim is to illustrate how the Adaptive Buffer Pool effectively reduces memory usage in different situations. The comparison is shown in Figure 11.
In these configurations, the Adaptive Buffer Pool consistently reduces memory usage by adjusting to the specific tensor sizes of each model, achieving an average reduction of 72.71% in different models. An interesting finding is that, for the Qwen2.5-14B and Qwen2.5-32B models, the original SSD offloading system uses the same amount of memory, while the Adaptive Buffer Pool shows slightly higher memory usage for the Qwen2.5-32B model. This occurs because the largest layer, the embedding layer, is identical in both the Qwen2.5-14B and Qwen2.5-32B models, leading to equal memory usage in the original system, which is limited by the largest layer size. However, in the Qwen2.5-32B model, the feedforward-related weights are larger, resulting in slightly increased memory usage compared to the Qwen2.5-14B. This finding highlights that the Adaptive Buffer Pool effectively reveals the actual memory needs of different models.
Fused Overflow Check Kernel
To assess the impact of the method described in Section IV-D, the analysis focuses on latency as a performance metric and memory efficiency compared to the implementation of the original SSD system. Figure 10(a) shows the overflow check latency across various models on Server Configuration 1, which uses an Intel(R) Xeon(R) 6780E CPU. In contrast, Figure 10(b) shows the corresponding results for Server Configuration 2, equipped with dual AMD EPYC 7282 CPUs. The latency scales linearly with model size, as the operation examines each gradient for overflow. The Fused Overflow Check Kernel, by removing intermediate tensor creation and multi-step processes, performs a single pass through the values, resulting in an average latency reduction of 97%. This kernel runs in every iteration that requires an update, with computational complexity scaling linearly with model size, underscoring the importance of performance optimization. For memory efficiency, Figure 10 indicates that MemAscend does not have additional memory overhead compared to ZeRO-Infinity. Unlike the original overflow check, which generates multiple intermediate values and incurs significant memory overhead, the fused kernel eliminates this overhead by avoiding intermediate computations entirely.
@@
Figure 12: Comparison of read and write latency and bandwidth between ZeRO-Infinity and MemAscend across various tensor sizes.
This evaluation examines the performance of the Direct NVMe Engine, described in Section IV-E, for SSD I/O in training scenarios with SSD offloading. The engine manages offloading of fp16 and fp32 model weights, gradients, and optimizer states. To determine suitable tensor sizes for benchmarking, model weight distributions were analyzed, excluding tensors too small to benefit from offloading. For instance, tensors with fewer than two million elements perform better in CPU memory. Common dense LLM architectures include several layers suitable for SSD offloading, such as feedforward layers, q, k, v, and o projection layers, the embedding layer, and the linear head layer. To ensure a precise comparison with the original SSD offloading system, tensor size ranges for fp16 and fp32 data were collected across various models. The analysis also considers multi-GPU setups, where data splits across processes. For example, a 100-million-element tensor divides into 50 million elements per process in a two-GPU configuration. Benchmarks used Configuration 2, as specified in Table II. The original SSD offloading system, reliant on filesystem I/O, utilized two SSDs in a RAID0 array formatted with ext4. By contrast, MemAscend employs the Direct NVMe Engine, which bypasses software RAID and sends requests directly to SSDs through the NVMe driver.
Figures 12(a) and 12(b) display write latency and bandwidth across tensor sizes. The Direct NVMe Engine consistently lowers write latency. For a 2,097,152-byte tensor, latency decreases from 988.222 microseconds to 219.418 microseconds, and for a 3,114,270,720-byte tensor, it drops from 304,604.779 to 266,212.637 microseconds. This improvement results from eliminating filesystem and RAID layer overhead. The engine also achieves higher write bandwidth across all tested sizes. As tensor sizes increase, the performance gap narrows, likely because filesystem and RAID overhead becomes less significant when host computation limits performance. Figures 12(c) and 12(d) show read latency and bandwidth. Read latency remains similar to the filesystem approach, as reads involve less overhead without costly allocation. However, the Direct NVMe Engine provides more consistent read bandwidth, while the filesystem-based method exhibits greater variability. The Direct NVMe Engine delivers superior write performance and more stable read throughput, surpassing the original SSD offloading system in SSD offloading training workflows.

### v2

@@ IV-B Adaptive Buffer Pool
IV-C Alignment Free Pinned Memory Allocation
@@
As noted in Section III-B, PyTorch’s standard caching allocator applies a power-of-two alignment policy for pinned memory allocations, which leads to significant memory waste, particularly when dealing with large static buffers used in SSD offloading. To mitigate this, a custom C++ extension has been included in the proposed Memasend. It allocates CPU memory directly using posix_memalign, aligning only to 4096 bytes to meet DMA requirements. After allocation, the memory is page-locked and registered with all CUDA contexts using the CUDA API cudaHostRegister, with the cudaHostRegisterPortable flag. A CPU-pinned tensor is then created via the torch::from_blob API, enabling efficient GPU DMA transfers. This direct allocation and registration approach ensures that pinned memory buffers occupy only the required space plus minimal alignment, avoiding the internal fragmentation caused by power-of-two rounding. As a result, system memory usage drops significantly for large allocations such as the gradient partition flat buffer and the adaptive buffer pool.
IV-D Fused Overflow Check Mechanism
According to the discussion in Section III-C, the SSD offloading system detects gradient overflows by invoking operations, such as isnan() and isinf(), and leads to peak memory usage when transitioning between these functions. Combining these checks into a single pass removes the intermediate tensors each call would otherwise create, lowering peak memory demand. Gradients are produced in fp16 during back-propagation, yet the partition’s flat buffer is retained in fp32 for accumulation, so fp16-level overflows must still be vetted against fp32 representations. Algorithm 1 illustrates the fused procedure: under IEEE 754, overflow manifests when all exponent bits equal 1; NaN values (Not a Number) add a non-zero mantissa, whereas positive or negative infinity has a zero mantissa. Line 2 defines this bit pattern, and lines 5–8 apply bitwise logic to flag any value with an all-ones exponent as overflowed. Each element is independent, enabling wide parallel execution; practical tests show roughly 97% efficiency for models spanning 1 B–32 B parameters. By preventing intermediate allocations and exploiting bit-level checks, the fused mechanism removes the overflow-detection bottleneck at scale.
Algorithm 1 Fused Overflow Check
1:Gradient tensor $G[0\,\dots\,N-1]$
2:Boolean overflow
3:overflow $\leftarrow$ False
4:EXP_ALL_ONES_MASK $\leftarrow$ 0x7F800000
5:for each $i=0$ to $N-1$ in parallel do
6:     $bits\leftarrow\textsc{reinterpret\_uint32}(G[i])$
[…]
8:         overflow $\leftarrow$ True
9:         break $\triangleright$ Early exit from all threads
10:     end if
11:end for
12:return overflow
Figure 7: Workflow of the direct NVMe engine for optimized SSD I/O.
@@ IV-E Direct NVMe Engine
The primary goal of the direct NVMe engine is to bypass the Linux filesystem overhead by using raw I/O to directly send asynchronous I/O (AIO) requests to the device driver via iouring or libaio. The overview of architecture and its workflow is shown in Figure 7. First, the SSD offloading system issues a tensor AIO request. The request context contains a unique tensor name (as a key), a data pointer, and an offset aligned to the tensor size. Next, the direct NVMe engine distributes this request among multiple thread workers, dividing the data into equal portions for balanced performance. Each thread consults the tensor location dictionary to determine which SSD device and logical block address (LBA) to use. If the tensor location dictionary has not recorded the tensor yet, it consults the location allocator for SSD space. This step ensures that the data are horizontally partitioned across SSDs, effectively implementing striping in place of Software RAID 0. The shared device information structure, implemented via shared memory, guarantees that multiple processes do not allocate overlapping SSD regions. After each thread obtains its data pointer, offset, LBA, and raw SSD device path, it sends a raw AIO request through libaio. By applying load balancing, stripping, and multi-threading, the Direct NVMe Engine reduces latency and achieves higher throughput compared to approaches that rely on filesystem-based AIO.
@@ V Memory Efficiency and Performance Analysis
By streamlining memory consumption, the MemAscend framework reduces system memory demand relative to earlier SSD offloading techniques. Those lower system memory footprints can translate directly into cheaper training setups; in other words, the same hardware budget can now accommodate models with more trainable parameters. The lower system memory usage also unlocks longer context windows, because the larger activation-checkpoint tensors required for extended sequences can be moved from GPUs to system memory without exceeding capacity. Therefore, instead of slowing down the training process, the overall performance benefits follow as well. Meanwhile, the extra system memory headroom permits larger batch sizes, driving GPUs toward higher utilization and boosting overall throughput. The upcoming analysis examines these gains in detail and explains how the reclaimed RAM supports longer context lengths and bigger batches, both of which contribute to faster training at lower cost. These experiments are conducted using Configuration 1, as described in Table II.
@@ V-A Memory Efficiency
Figure 10 first breaks down system memory usage while training the Qwen2.5-7B model [34], highlighting substantial savings in every major component. Notably, both the compared Zero-Infinity and the proposed MemAscend are integrated with the Liger-Kernel, Flash-Attention, and offloaded gradient checkpointing for reducing residual memory usage. In the proposed MemAscend, the adaptive buffer pool is first introduced to reduce memory fragmentation. For the Qwen2.5-7B model in fp16, this approach requires about 2.46 GiB of memory for buffer allocations. In contrast, the standard SSD offloading method reserves a fixed 9.14 GiB. This change cuts memory use in the parameter buffer pool by 73%. To eliminate peak memory spikes from overflow checking, MemAscend implements the fused overflow check mechanism. Conventional checks rely on intermediate buffers, adding a 1.25$\times$ overhead and consuming 35.46 GiB. The fused mechanism maintains numerical stability without this overhead, reducing usage to 0 GiB. SSD offloading also depends on pinned memory to move data between CPU and GPU. Traditional allocation uses power-of-two alignment, introducing significant waste. MemAscend applies the alignment-free pinned memory allocation that allocates only the required memory. This reduces pinned memory overhead from 24.90 GiB to 1.63 GiB, a 93.5% decrease. Some components remain constant across methods, including the gradient flat buffer and small system allocations like optimizer state buffers and the swap-out buffer. These account for stable memory usage of roughly 28.37 GiB and 11.17 GiB, respectively. In summary, MemAscend brings total peak system memory usage down from 109.04 GiB under ZeRO-Infinity to 43.64 GiB, a 60% reduction. This allows for training larger models and handling longer input sequences within the same memory budget.
V-B Enabling Larger Context Length
In the SSD offloading scenario, the key constraint in expanding the context length or batch size is the activation checkpoint value, which scales linearly with both parameters. Within the SSD offloading framework, to alleviate this GPU memory bottleneck, a pinned memory buffer on the CPU can be utilized for swapping activation checkpoint values from GPU memory to system memory, enabling longer context lengths. However, this shifts the bottleneck from GPU memory to system memory. Notably, the total size of activation checkpointed values in system memory is calculated using the following formula, where $N_{g}$ denotes the GPU count, $B$ the batch size, $C$ the context length, $L$ the number of layers, $H$ the hidden size, $F_{16}$ as the bytes per fp16, and $P_{m}$ the pinned memory overhead. Here, the number of layers $L$ accounts for the activation checkpointing enabled in each transformer layer.
@@ V-B Enabling Larger Context Length
Figure 10 shows that the usage of system memory increases with longer context length when running Qwen2.5-7B on two GPUs. Our MemAscend method scales linearly with context length, exhibiting minimal pinned memory overhead, meaning the increase is almost entirely due to context length scaling. In contrast, the original SSD offloading system scales more rapidly and begins with higher memory usage. This disparity arises because its pinned overhead aligns to the nearest power-of-two value, resulting in faster scaling and instances where different context lengths yield identical memory usage due to this alignment. Consequently, our method significantly reduces system memory consumption and slows the scaling rate with increasing context length, offering greater flexibility to extend context length without exhausting system memory. For example, with a 128 GiB system memory limit, the original SSD offloading system supports a context length of 16,384, while MemAscend scales up to 131,072 under the same configuration.
V-C Enabling Higher Throughput
The proposed MemAscend not only supports a larger trainable context length but also unlocks the potential for higher throughput (tokens/sec) by accommodating larger batch sizes through reclaimed system memory budget. Larger batch sizes allow GPUs to process more tokens under the same amount of model weights transfers, thus improving the compute-to-transfer ratio and offsetting SSD offloading overhead. Figure 10 compares the baseline SSD offloading technique with MemAscend during Qwen2.5-7B training on two GPUs. Memory-usage trends remain proportional to batch size or context length, yet MemAscend’s curve rises far more gently, leaving headroom to keep increasing the batch without exhausting host RAM. Consequently, the throughput of MemAscend scales almost linearly with batch size, confirming that more samples per offload cycle translate into higher overall performance. For instance, under a 128 GiB system-memory cap, the baseline tops out at batch 4, whereas MemAscend reaches 32, converting the reclaimed RAM into an eight-fold boost in tokens per second.
VI Experimental Evaluation
VI-A Setup
All experiments were carried out on two servers. Detailed hardware specifications are provided in Table II. MemAscend is evaluated using mixed-precision training of large language models with SSD offloading. The workloads vary by model architecture, context length, and batch size. Models include Llama3.1-8B and Qwen2.5, with parameter sizes of 7 billion, 14 billion and 32 billion. All models use the Hugging Face Transformers library (v4.46.3). The context lengths range from 4,096 to 131,072 tokens. Batch sizes vary across different experiments. The main evaluation metric is peak system memory usage during training and training throughput in tokens per second to assess the overall performance.
@@ VI-A Setup
To evaluate the effectiveness of MemAscend, the SOTA ZeRO-Infinity is compared, and both systems share a common setup to ensure fairness. Static memory components, including model parameters and optimizer states, are stored entirely on the SSD. To reduce GPU memory usage, gradient checkpointing, Flash-Attention, and the Liger-Kernel are enabled for both setups. For experiments that require offloaded gradient checkpointing, a modified version of Unsloth is utilized, adapted from its original single-GPU design.
@@ VI-B1 Assessment of Individual Methods
As each model varies in weight tensor size due to differences in vocabulary size, hidden size, and other factors, this experimental section presents the memory usage of the buffer pool across various models. The aim is to illustrate the effectiveness of the proposed adaptive buffer pool in reducing memory usage in different situations. The comparison is shown in Figure 11. In these configurations, the adaptive buffer pool consistently reduces memory usage by adjusting to the specific tensor sizes of each model, achieving an average reduction of 72.71% accross different models. An interesting finding is that, for the Qwen2.5-14B and Qwen2.5-32B models, the original SSD offloading system uses the same amount of memory, while the adaptive buffer pool shows slightly higher memory usage for the Qwen2.5-32B model. This occurs because the largest layer, the embedding layer, is identical in both the Qwen2.5-14B and Qwen2.5-32B models, leading to equal memory usage in the original system, which is limited by the largest layer size. However, in the Qwen2.5-32B model, the feedforward-related weights are larger, resulting in slightly increased memory usage compared to the Qwen2.5-14B. This finding highlights that the adaptive buffer pool effectively reveals the actual memory needs of different models.
Fused Overflow Check Mechanism
To assess the impact of the method described in Section IV-D, the analysis focuses on latency as a performance metric and memory efficiency compared to the implementation of the original SSD system. Figure 13(a) shows the overflow check latency across various models on Server Configuration 1, which uses an Intel(R) Xeon(R) 6780E CPU. In contrast, Figure 13(b) shows the corresponding results for Server Configuration 2, equipped with dual AMD EPYC 7282 CPUs. The latency scales linearly with model size, as the operation examines each gradient for overflow. The fused overflow check mechanism, by removing intermediate tensor creation and multi-step processes, performs a single pass through the values, resulting in an average latency reduction of 97%. This mechanism runs in every iteration that requires an update, with computational complexity scaling linearly with model size, underscoring the importance of performance optimization. For memory efficiency, Figure 13 indicates that MemAscend does not have additional memory overhead compared to ZeRO-Infinity. Unlike the original overflow check, which generates multiple intermediate values and incurs significant memory overhead, the fused mechanism eliminates this overhead by avoiding intermediate computations.
@@
Figure 14: Comparison of NVMe SSD read/write latency and bandwidth between the ZeRO-Infinity approach and the direct NVMe engine across various tensor sizes.
This evaluation examines the performance of the direct NVMe engine, described in Section IV-E, for SSD I/O in training scenarios with SSD offloading. The engine manages offloading of fp16 and fp32 model weights, gradients, and optimizer states. To determine suitable tensor sizes for benchmarking, model weight distributions were analyzed, excluding tensors that are too small to benefit from offloading. For instance, tensors with fewer than two million elements perform better in CPU memory. Common dense LLM architectures include several layers suitable for SSD offloading, such as feedforward layers, q, k, v, and o projection layers, the embedding layer, and the linear head layer. To ensure a precise comparison with the original SSD offloading technique, tensor size ranges for fp16 and fp32 data were collected across various models. The analysis also considers multi-GPU setups, where data are split across processes. For example, a 100-million-element tensor divides into 50 million elements per process in a two-GPU configuration. Benchmarks used Configuration 2, as specified in Table II. The original SSD offloading system, reliant on filesystem I/O, utilized two SSDs in a RAID0 array formatted with ext4. By contrast, MemAscend employs the eirect NVMe engine, which bypasses software RAID and sends requests directly to SSDs through the NVMe driver.
Figures 14(a) and 14(b) display write latency and bandwidth across tensor sizes. The direct NVMe engine consistently lowers write latency. For a 2,097,152-byte tensor, latency decreases from 988.222 microseconds to 219.418 microseconds, and for a 3,114,270,720-byte tensor, it drops from 304,604.779 to 266,212.637 microseconds. This improvement results from eliminating filesystem and RAID layer overhead. The engine also delivers higher write bandwidth across all tested sizes, achieving up to 4.5x gains and an average improvement of 72.04%. As tensor sizes increase, the performance gap narrows, likely because filesystem and RAID overhead become less significant when host computation dominates performance. Figures 14(c) and 14(d) show read latency and bandwidth. Read latency remains similar to the filesystem approach, as reads involve less overhead without costly allocation. However, the direct NVMe engine provides more consistent read bandwidth, while the filesystem-based method exhibits greater variability. The direct NVMe engine delivers superior write performance and more stable read throughput, surpassing the original SSD offloading technique.